Baba Company is a manufacturing firm that uses job-order costing. The company's inventory balances were as...

  1. Baba Company is a manufacturing firm that uses job-order costing. The company's inventory balances were as follows at the beginning and end of the year:

           

Beginning Balance

Ending Balance

Raw materials....................

$22,000

$25,000

Work in process.................

$52,000

$34,000

Finished goods..................

$121,000

$136,000

The company applies overhead to jobs using a predetermined overhead rate based on machine-hours. At the beginning of the year, the company estimated that it would work 40,000 machine-hours and incur $200,000 in manufacturing overhead cost. The following transactions were recorded for the year:

  • Raw materials were purchased, $412,000.
  • Raw materials were requisitioned for use in production, $409,000 $(342,000 direct and $67,000 indirect).
  • The following employee costs were incurred: direct labor, $314,000; indirect labor, $67,000; and administrative salaries, $129,000.
  • Selling costs, $135,000.
  • Factory utility costs, $25,000.
  • Depreciation for the year was $102,000 of which $90,000 is related to factory operations and $12,000 is related to selling, general, and administrative activities.
  • Manufacturing overhead was applied to jobs. The actual level of activity for the year was 44,000 machine-hours.
  • Sales for the year totaled $1,198,000.

           

            Required:

  1. What is the Company’s Predetermined Overhead Rate
  2. What is the Amount of Allocated Overhead
  3. What is the amount of Actual Overhead
  4. Was the overhead under-allocated or over-allocated? By how much?

Prepare the Journal Entry to close out the over/under allocated
Answer #1

a) Predetermine overhead rate = 200000/40000 = $5 per machine hour

b) Amount of allocated overhead = 44000*5 = $220000

c) Actual overhead = 67000+67000+25000+90000 = 249000

d) Under allocated overhead = 249000-220000 = $29000

Journal entry

Date account and explanation debit credit
Cost of goods sold 29000
Manufacturing overhead 29000
(To record under allocated overhead)